A new car is purchased for 24500 dollars. The value of the car depreciates at 14.75% per year. What will the value of the car be, to the nearest cent, after 14 years?


1
Expert's answer
2022-05-23T14:01:51-0400

We would apply the formula for exponential decay which is expressed as

"A=P(1-r)^t"

Where

A - represents the value of the car after t years.

t - represents the number of years.

P - represents the value of the car.

r - represents rate of decay.

From the information given,

P = 24500 dollars

r = 14.75% = 14.75/100 = 0.1475

t = 14 years

Therefore

"A=24500(1-0.1475)^{14}=24500(0.8525)^{14}=2623.54"